OpenStudy (anonymous):

solve 1/x - 5/2x^2 = 8/5x

OpenStudy (whpalmer4):

You really need to either use the equation editor or sufficient parentheses to make your meaning unambiguous. Is the middle term \[\frac{5}{2x^2}\]or is it\[\frac{5}{2}x^2\]Makes a big difference in the answer! I'm going to assume this is what you meant: \[\frac{1}{x}-\frac{5}{2x^2}=\frac{8}{5x}\] My inclination would be to get rid of the fractions by multiplying through by all of the denominators: \[x*2x^2*5x(\frac{1}{x}-\frac{5}{2x^2}) = x*2x^2*5x*\frac{8}{5x}\]\[10x^3-25x^2=16x^3\]Now factor out the common factors \[x^2(10x-25) = x^2*16x\]Divide through by the common factor\[10x-25=16x\]I trust you can see how to solve it from here!

OpenStudy (anonymous):

wait but there is no common factor :(

OpenStudy (anonymous):

by the way you interpreted the question correctly, sorry for the confusion

OpenStudy (whpalmer4):

the common factor was \(x^2\)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I ended up with 10x^2-8 as my answer

OpenStudy (anonymous):

but i don't think thats right

OpenStudy (whpalmer4):

\[\color\red{ x^2}(10x-25) = \color\red{ x^2}(16x)\]

OpenStudy (anonymous):

but where did the x^2 come from?

OpenStudy (anonymous):

nevermind :) sorry

OpenStudy (whpalmer4):

I factored it out: \[10x^3-25x^2=16x^3\]\[x^2*10x-x^2*25=x^2*16x\]\[x^2(10x-25) = x^2(16x)\] \[10x-25=16x\]Subtract \(10x\) from each side\[-25=6x\]\[x=-\frac{25}{6}\]

OpenStudy (anonymous):

so did u get -4.12?

OpenStudy (anonymous):

x = -4.12

OpenStudy (anonymous):

yes ok or -25/6... ok, thanks :)

OpenStudy (whpalmer4):

No, you should leave it as \[x=-\frac{25}{6}\]unless you've been specifically asked for a decimal number. And that decimal number would be -4.16666666666666666666666 etc.

OpenStudy (whpalmer4):

why would you want an inconvenient, only approximate answer when you could have a compact, exact answer? :-)

OpenStudy (anonymous):

I don't know :)

OpenStudy (anonymous):

what about 2x-3/x + 4 = x - 3/x ? I got x^2 - 6x

OpenStudy (anonymous):

I tried to do one of those equations so you could read it better but I was having trouble using it...

OpenStudy (whpalmer4):

Okay, then put parentheses around each numerator or denominator or fraction the original problem could have been written (1/x) - (5/(2x^2)) = 8/(5x) and there would be no question how to interpret it

OpenStudy (anonymous):

(2x-3/x) + 4 = x - (3/x)

OpenStudy (whpalmer4):

Better, but I still can't tell if the first term is \[\frac{2x-3}{x}\] or\[2x-\frac{3}{x}\]